### v1

# Optimal observables for (non-)equilibrium quantum metrology from the master equation

Abstract
We demonstrate how observables with optimal sensitivity to environmental properties can be constructed explicitly from the master equation of an open-quantum system. Our approach does not rely on the explicit solution of the master equation. This makes the symmetric logarithmic derivative (SLD), the operator of optimal sensitivity and key quantity in quantum metrology, available to a large class of systems of interest, both in and out-of-equilibrium. We validate our approach by reproducing the SLD for temperature in quantum Brownian motion and demonstrate its versatility by constructing the optimal observable for the non-equilibrium relaxation rate.
preprint: APS/123-QED
I Introduction
Quantum metrology exploits the quantum properties of systems to estimate physical parameters with the highest possible precision, potentially surpassing classical limits (for a recent review see [1]). A central object in local estimation theory [2] is the symmetric logarithmic derivative (SLD), $\hat{\Lambda}_{\theta}$ [3, 4], which determines the quantum Fisher information (QFI), a measure of the ultimate sensitivity with which a parameter $\theta$ encoded in a quantum state $\hat{\rho}(\theta)$ can be estimated. Defined implicitly via the generalized Lyapunov equation
$\partial_{\theta}\hat{\rho}(\theta)=\frac{1}{2}\left(\hat{\Lambda}_{\theta}\hat{\rho}(\theta)+\hat{\rho}(\theta)\hat{\Lambda}_{\theta}\right)-\hat{\rho}(\theta)\langle\hat{\Lambda}_{\theta}\rangle,$ (1)
the SLD plays a dual role: it determines both the QFI as $\mathcal{F}_{\theta}=\operatorname{Tr}[\hat{\rho}(\theta)\hat{\Lambda}_{\theta}^{2}]$, and more fundamentally, the optimal observable whose variance bounds the achievable estimation error in the quantum Cramér-Rao inequality [5].
Despite its central conceptual role, the practical use of the SLD in quantum metrology remains constrained [1]. Its computation is often contingent on prior knowledge of the explicit solution $\hat{\rho}(\theta)$ to the dynamical equation governing the quantum system, readily available predominantly in Gaussian states [6] or two-level systems. This limitation is particularly restrictive in open quantum systems [7], used for impurity metrology, where the explicit density matrix is typically available only through numerical simulations or perturbative expansions, such as in the study of heavy quark impurities in high energy nuclear physics [8, 9, 10, 11].
In the context of equilibrium thermometry, for instance, it is common to evaluate the QFI for thermal (Gibbs) states by exploiting analytic forms of $\hat{\rho}(T)$. Formally the SLD may be expressed as a simple function of the system Hamiltonian, such as $\hat{\Lambda}_{T}=(\hat{H}-\langle H\rangle)/T^{2}$ [12] but direct energy measurements may not be feasible in practice. An interesting example is Ref. [13] where the SLD for a heavy bosonic impurity was explicitly constructed
$\hat{\Lambda}_{T}$ $=C_{x}\left(\hat{x}^{2}-\langle{x}^{2}\rangle\right)+C_{p}\left(\hat{p}^{2}-\langle{p}^{2}\rangle\right)$ (2)
making reference to the expectation values of practically accessible operators, such as position $\hat{x}$ and momentum $\hat{p}$ in the coefficients
$C_{x}$ $=\frac{4\langle{p}^{2}\rangle^{2}\,\chi_{T}(\hat{x}^{2})+\hbar^{2}\chi_{T}(\hat{p}^{2})}{8\langle{x}^{2}\rangle^{2}\langle{p}^{2}\rangle^{2}-\hbar^{4}/2},$ (3)
and $C_{p}$ for which one replaces $\hat{x}\leftrightarrow\hat{p}$ in $C_{x}$. Note however that in the above expression the temperature susceptibilities $\chi_{T}(\hat{x}^{2})$ and $\chi_{T}(\hat{x}^{2})$ enter, which require a microscopic model and access to the explicit density matrix to produce predictions for $\chi_{\theta}(\hat{\cal O})=\partial_{\theta}{\rm Tr}[\hat{\rho}(\theta)\hat{\cal O}]$.
In non-equilibrium thermometry, the SLD has been investigated for various two-level (qubit) systems. Let us consider the interesting example of in-situ thermometry with dephasing impurities pioneered in Ref. [14]. In this scenario, due to exact energy conservation of the probe, thermalization with the surrounding environment is suppressed. Even in the absence of thermalization the probe can develop correlations with the environment that give access to temperature and an SLD has been constructed
$\hat{\Lambda}_{T}$ $\propto\cos(\varphi)\,\hat{\sigma}_{\parallel}+\sin(\varphi)\,\hat{\sigma}_{\perp},$ (4)
$\tan(\varphi)$ $=\frac{|v|(1-|v|)^{2}\,\partial_{T}\phi}{\partial_{T}|v|}.$ (5)
Here $|v|$ and $\phi$ parameterize the density matrix of the probe and $\hat{\sigma}_{\parallel}$ and $\hat{\sigma}_{\perp}$ denote linear combinations of spin measurement operators in the $x$ and $y$ direction. Again we see that derivatives of the explicit solution of the density matrix are required for predictions, in the form of $\partial_{T}|v|$ and $\partial_{T}\phi$ analytically known only in weakly-coupled environments.
In this work our goal is to facilitate theory predictions and to elevate the utility of the SLD to a larger class of systems, by developing a novel method to construct and evaluate the SLD directly from the master equation, without requiring the explicit form of the density matrix $\hat{\rho}(\theta)$. Our approach leverages the structure of the open-system generator, whether time-local or non-Markovian, and manage to systematically propagate the SLD alongside the dynamical evolution. This formulation makes it possible to identify optimal observables at arbitrary times, in particular in the transient regime, where the system has not reached a steady state and the full density matrix is often not analytically tractable.
By disentangling the determination of the SLD from the solution of $\hat{\rho}(t)$, our method enables metrological optimization in a broad class of experimentally relevant situations, including driven systems, nonequilibrium dynamics, and strongly coupled probes. As such, it provides a valubale tool in the theoretical arsenal of quantum metrology.
This paper is structured as follows. In section II we introduce our formalism to determine the SLD from the master equation. Section III applies the formalism to thermometry in the equilibrium steady state of quantum Brownian motion with a harmonic potential, before constructing in section IV the SLD for both temperature and relaxation rate in a non-equilibrium scenario based on a squeezed Gaussian state.
II Formalism
In this section, we present a method for the explicit construction of the symmetric logarithmic derivative (SLD) $\hat{\Lambda}_{\theta}$, corresponding to a given environmental parameter $\theta$.
We begin by considering a general master equation governing the evolution of the reduced density matrix $\rho_{S}(t,\theta)$ of an open quantum system,
$\partial_{t}\hat{\rho}_{S}(t,\theta)=\hat{\mathcal{L}}[\theta]\hat{\rho}_{S}(t,\theta),$ (6)
where $\hat{\mathcal{L}}$ denotes a linear superoperator, which may depend explicitly on the environmental parameter $\theta$. Our formulation does not require the dynamics to be Markovian or weakly coupled, only that the master equation is linear in $\rho_{S}$.
The central idea of our approach is to exploit Schwarz’s theorem on the symmetry of mixed partial derivatives for twice differentiable functions. This allows us to relate the temporal derivative in the master equation (6) to the parameter derivative appearing in the definition of the SLD via
$\partial_{\theta}\left(\partial_{t}\hat{\rho}(t,\theta)\right)=\partial_{t}\left(\partial_{\theta}\hat{\rho}(t,\theta)\right).$ (7)
This identity serves as the bridge between the dynamics and the sensitivity operator.
For Markovian evolution, it is known that differentiability of the initial state guarantees differentiability of the evolved state [15]. In general, twice differentiability requires additional smoothness conditions on the generator $\mathcal{L}$, which must be verified on a case-by-case basis. For instance, an explicit analytic treatment of the Caldeira-Leggett master equation [16] confirms that twice differentiability holds for quantum Brownian motion of a point particle in an Ohmic environment [7].
To proceed, we adopt a time-independent operator expansion for the SLD,
$\hat{\Lambda}_{\theta}=\sum_{i}c^{(i)}_{\theta}\hat{A}_{i},$ (8)
where the $\hat{A}_{i}$ are a fixed set of Hermitian operators and the coefficients $c^{(i)}_{\theta}$ encode the parameter sensitivity. While the SLD is independent of time, its expectation value may acquire time dependence through the dynamical evolution of the operators $\langle\hat{A}_{i}\rangle(t)$.
Inserting the ansatz (8) into the definition of the SLD and evaluating Eq. (7) at the level of expectation values yields a system of equations for the unknown coefficients $c^{(i)}_{\theta}$. Specifically, multiplying both sides of Eq. (7) by a test operator $\hat{A}_{j}$ and taking the trace leads to
$\mathrm{Tr}\left[\partial_{\theta}\partial_{t}\hat{\rho}(t,\theta)\,\hat{A}_{j}\right]=\mathrm{Tr}\left[\partial_{t}\partial_{\theta}\hat{\rho}(t,\theta)\,\hat{A}_{j}\right].$ (9)
Upon substituting the SLD expansion (8) and computing the partial derivatives, the resulting expression decomposes into two terms: one that is independent of the coefficients $c^{(i)}_{\theta}$ and linear in $\hat{A}_{j}$, denoted $D_{j}$, and another that depends linearly on $c^{(i)}_{\theta}$ and involves operator bilinears, denoted $M_{ji}$. This yields a linear system,
$\sum_{i}M_{ji}\,c^{(i)}_{\theta}=D_{j},$ (10)
which determines the coefficients of the optimal observable. Provided that the matrix $M_{ji}$ is invertible – a condition that depends on the choice of basis operators and the structure of the master equation – the SLD can be constructed explicitly in terms of dynamical expectation values.
III Equilibrium probe thermometry in Quantum Brownian Motion
In this section we verify our formalism by constructing the optimally sensitive observable for temperature measurements in the steady state of quantum Ornstein-Uhlenbeck motion [17, 18], i.e. quantum Brownian motion in the presence of a harmonic potential [7]. This motion describes the dynamics of a confined and heavy point particle impurity in an Ohmic environment and leads to a stationary state at late time with time independent expectation values. We use the Caldeira-Leggett (CL) master equation [19, 20] in its original, non-Lindblad form to follow the evolution of the underlying reduced density matrix of the impurity. The key result is that our approach reproduces the known solution for the SLD, obtained from the explicit solution of the master equation in the late-time limit.
Making use solely of the master equation, let us inspect it in more detail. Denoting the single particle Hamiltonian by $\hat{H}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega^{2}\hat{x}^{2}=b\hat{p}^{2}+c\hat{x}^{2}$, we have
$\partial_{t}\hat{\rho}(t,T)=-i[\hat{H},\hat{\rho}]-i\gamma[\hat{x},\{\hat{p},\hat{\rho}\}]-2\gamma mT[\hat{x},[\hat{x},\hat{\rho}]],$ (11)
[…]
In the absence of inner structure the two hermitian operators that characterize all motion of the probe are the position $\hat{x}$ and momentum $\hat{p}$ operator. We thus consider powers of these two operators as basis for the operator expansion of eq. 8. Since one and two-point functions are the most easily accessible quantities in experiment, our truncated basis reads explicitly
$\hat{A}_{j}\in\{\hat{x},\hat{p},\hat{x}^{2},\hat{p}^{2},\{\hat{x},\hat{p}\}\},$ (12)
where the last term denotes the anticommutator of position and momentum.
They key relation eq. 7 requires us to take the derivative w.r.t. the environment parameter of interest, here $T$, of the master equation. Whenever $\partial_{T}$ is applied to the density matrix, $\partial_{T}\hat{\rho}$ will be replaced by the definition of the SLD eq. 1. Temperature appears explicitly on the r.h.s. of eq. 11. In turn after applying $\partial_{T}$, there remains one single terms where no temperature derivative of the density matrix occurs. After converting the operator expressions to expectation values with the help of a test operator $\hat{A}_{j}$ via eq. 9 this single term will become $D_{j}$ in eq. 10. The remaining terms depend linearly on $\Lambda$, so that when the SLD too is expanded in an operator basis via eq. 8, they furnish the l.h.s. of eq. 10.
The linear system of equations that defines the expansion coefficients $c^{(i)}_{T}$ of the SLD thus reads
$M_{ij}=$ (13)
$-\tfrac{1}{2}\left\langle\left\{\hat{A}_{i}\left[\hat{H},\hat{A}_{j}\right]\right\}\right\rangle-\tfrac{i}{2}\left\langle\left\{\hat{A}_{i},\left\{\hat{p},\left[\hat{A}_{j},\hat{x}\right]\right\}\right\}\right\rangle$
$-\tfrac{\gamma T}{2b}\left\langle\left\{\hat{A}_{i},\left[\hat{x},\left[\hat{x},\hat{A}_{j}\right]\right]\right\}\right\rangle+\tfrac{i\gamma}{2}\left\langle\left\{\hat{p},\left[\left\{\hat{A}_{i},\hat{A}_{j}\right\},\hat{x}\right]\right\}\right\rangle$
$+\tfrac{\gamma T}{2b}\left\langle\left[\hat{x},\left[\hat{x},\left\{\hat{A}_{i},\hat{A}_{j}\right\}\right]\right]\right\rangle-i\left\langle\hat{A}_{j}\right\rangle\left\langle\left[\hat{A}_{i},\hat{H}\right]\right\rangle$
$-i\gamma\left\langle\hat{A}_{j}\right\rangle\left\langle\left\{\hat{p},\left[\hat{A}_{i},\hat{x}\right]\right\}\right\rangle-\tfrac{\gamma T}{b}\left\langle\hat{A}_{j}\right\rangle\left\langle\left[\hat{x},\left[\hat{x},\hat{A}_{i}\right]\right]\right\rangle$
with the constant terms
$D_{j}=\frac{\gamma}{b}\left\langle\left[\hat{x},\left[\hat{x},\hat{A}_{j}\right]\right]\right\rangle.$ (14)
Let us evaluate the entries of the matrix $M$ and the vector $D$ in the late-time steady state, where the system has become Gaussian. For an efficient handling of terms with different powers of the $\hat{x}$ and $\hat{p}$ operator we collect them in terms of the Weyl symbols [21]
$W(\hat{x}^{n}\hat{p}^{m})=\frac{1}{2^{n}}\sum_{k=0}^{n}\binom{n}{k}\hat{x}^{n-k}\hat{p}^{m}\hat{x}^{k}.$ (15)
The expectation value of any Weyl symbol with an odd power of either operator vanishes, such as e.g.
$\left\langle W(xp)\right\rangle=\frac{1}{2}\left\langle\left\{\hat{x},\hat{p}\right\}\right\rangle=0,$ (16)
$\left\langle W(x^{3}p)\right\rangle=0,\quad\left\langle W(xp^{3})\right\rangle=0,$ (17)
while the only relevant non-vanishing Weyl symbol is
$\left\langle W(x^{2}p^{2})\right\rangle=\langle x^{2}\rangle\langle p^{2}\rangle.$ (18)
Expectation values of higher powers of the individual operators in the Gaussian state factor in products of lower powers according to Wick’s theorem [22]. The relevant expressions are
$\langle x^{4}\rangle=3\langle x^{2}\rangle^{2},\quad\langle\hat{p}^{4}\rangle=3\langle\hat{p}^{2}\rangle$ (19)
$\langle x^{2}\rangle=\frac{a^{2}b}{4c},\quad\langle\hat{p}^{2}\rangle=\frac{a^{2}}{4},$ (20)
where we have introduced the abbreviation $a=\sqrt{4mT}$.
A careful evaluation of all $5\times 5$ entries of $M$ yields
$M=$ (21)
$\left[\begin{array}[]{ccccc}0&-2b\left\langle p^{2}\right\rangle&0&0&0\\[10.0pt] 2c\left\langle x^{2}\right\rangle&\begin{array}[]{c}2\gamma\left\langle p^{2}\right\rangle\\ -\frac{2\gamma T}{b}\end{array}&0&0&0\\[4.0pt] 0&0&0&2\gamma&-8b\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle-2b\\[4.0pt] 0&0&0&\begin{array}[]{c}8\gamma\left\langle p^{2}\right\rangle^{2}\\ -\frac{8\gamma T\left\langle p^{2}\right\rangle}{b}\end{array}&8c\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+2c\\[4.0pt] 0&0&\begin{array}[]{c}2b\\ +8c\left\langle x^{2}\right\rangle^{2}\end{array}&\begin{array}[]{c}-8b\left\langle p^{2}\right\rangle^{2}\\ -2c\end{array}&\begin{array}[]{c}-\frac{8\gamma T\left\langle x^{2}\right\rangle}{b}-2\gamma\\ +8\gamma\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle\end{array}\\ \end{array}\right]=\begin{bmatrix}0&-\frac{ba^{2}}{2}&0&0&0\\ \frac{a^{2}b}{2}&-\frac{\gamma a^{2}}{2}&0&0&0\\ 0&0&0&2\gamma&-\dfrac{a^{4}b^{2}}{2c}-2b\\ 0&0&0&-\dfrac{\gamma a^{4}}{2}&\frac{a^{4}b}{2}+2c\\ 0&0&\frac{a^{4}b^{2}}{2c}+2b&-\frac{a^{4}b}{2}-2c&-\frac{\gamma a^{4}b}{2c}-2\gamma\end{bmatrix},$ (37)
where after the second equal sign, the expectation values have been replaced by their explicit values in the late time limit, in terms of the model parameters. The constant term features a single non-vanishing entry and reads
$D=\left[0,0,0,-2\frac{\gamma}{b},0\right].$ (38)
The entries are ordered according to the ordering in eq. 12, i.e. the entry $M_{00}=M_{xx}$ and the entry $M_{44}=M_{\{x,p\}\{x,p\}}$.
Interesting structure emerges in the system of linear equations. Since the Caldeira-Leggett equation is not symmetric in position and momentum the matrix $M$ is not necessarily symmetric. In addition we find that the terms associated with a single power of the operators (left upper $2\times 2$ entries) factorize from the contributions related to products of operators (right bottom $3\times 3$ entries). Due to the fact that the constant terms vanish except for the entry $D_{p^{2}p^{2}}$, we obtain as solution
$c^{(x)}_{T}$ $=0,$ (39)
$c^{(p)}_{T}$ $=0,$
$c^{(x^{2})}_{T}$ $=\frac{4c\left(a^{4}b^{2}+4(bc+\gamma^{2})\right)}{a^{8}b^{4}-16b^{2}c^{2}},$
$c^{(p^{2})}_{T}$ $=\frac{4}{a^{4}b-4c},$
$c^{(\{x,p\})}_{T}$ $=\frac{16c\gamma}{a^{8}b^{3}-16bc^{2}}.$
Let us compare the expressions obtained here with the known non-vanishing solutions $\bar{c}^{(x^{2})}_{T}$ and $\bar{c}^{(p^{2})}_{T}$ obtained in Ref.[13] from the explicit Gaussian density matrix in the late time limit
$\bar{c}^{(x^{2})}_{T}$ $=\frac{4c\left(a^{4}b^{2}+4bc\right)}{a^{8}b^{4}-16b^{2}c^{2}},$ (40)
$\bar{c}^{(p^{2})}_{T}$ $=\frac{4}{a^{4}b-4c}.$ (41)
We find that in the Markovian approximation underlying the CL equation $\gamma\ll 2\pi T$ we can estimate $a^{4}b^{2}=4T^{2}\gg 4\gamma^{2}$ and thus our solution agrees with the known result.
Our approach however reveals an important detail. For the SLD to be well defined, i.e. to obtain a consistent solution of the linear system of equations the operator basis must include $\{\hat{x},\hat{p}\}$, even though the expectation value of that operator vanishes in the late time limit. The presence of $\gamma$ in our solution serves as a reminder that the system must be able to approach the late-time steady state, which is only possible with a finite value of $\gamma$, while the actual value of $\gamma$ does not play a role for the optimal observable itself
$\hat{\Lambda}_{T}=c^{(x^{2})}_{T}\hat{x}^{2}+c^{(p^{2})}_{T}\hat{p}^{2}.$ (42)
This result is encouraging, as we managed to construct an explicit expression for the SLD, based only on our knowledge of the master equation and which reproduces the known equilibrium result.
IV Non-equilibrium metrology for a squeezed gaussian state
In this section we apply our approach to non-equilibrium metrology, i.e. we consider a probe that is not yet in equilibrium with its environment. Our goal is to derive not only the optimally sensitive observable for the determination of the temperature of the environment, but also for the relaxation rate $\gamma$. The latter is not accessible in the late time steady state, where all information about the initial conditions, and thus the approach to equilibrium has been lost. As concrete realization we consider a squeezed Gaussian state in which $\langle\hat{x}\rangle=0$, $\langle\hat{p}\rangle=0$ but $\langle\{\hat{x},\hat{p}\}\rangle\neq 0$. Note that in general there does not exist a closed form explicit solution of the density matrix to the CL master equation in this regime and thus a direct evaluation of the SLD is not possible (for an analytic treatment of the CL master equation see e.g. [23, 16])
IV.1 Non-equilibrium Thermometry
For the non-equilibrium SLD for temperature we can simply reuse eq. 13 and eq. 14 but now have to evaluate the expectation values away from the steady state. This entails that mixed expectation values do not necessarily vanish but they factor according to Wick’s theorem [22], such as
$\langle W(x^{3}p)\rangle=\frac{3}{2}\langle x^{2}\rangle\langle\left\{\hat{x},\hat{p}\right\}\rangle,$ (43)
$\left\langle W(x^{2}p^{2})\right\rangle=\langle x^{2}\rangle\langle p^{2}\rangle+\frac{1}{2}\left\langle\left\{\hat{x},\hat{p}\right\}\right\rangle^{2}.$ (44)
Taking into account the additional terms contributing we obtain the following explicit expression for the matrix $M$:
Figure 1: Probe particle evolution in a squeezed Gaussian state at $\omega/m=4/3$ (top row) and $\omega/m=2/3$ (bottom row). Evolution of the expectation values (left) of the relevant observables $\langle x^{2}\rangle$ (green solid), $\langle p^{2}\rangle$ (teal solid), $\langle px+xp\rangle$ (purple solid). Their equilibrium values are shown as dashed gray lines. Corresponding expansion coefficients $c^{(i)}_{T}$ of the symmetric logarithmic derivative for temperature (right). The equilibrium solution of the coefficients is also shown as dashed gray lines.
$\begin{bmatrix}\begin{array}[]{c}-b\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}-2b\langle p^{2}\rangle\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}\\[20.0pt] \begin{array}[]{c}2c\langle x^{2}\rangle\\ +\gamma\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}2\gamma\langle p^{2}\rangle-\dfrac{2\gamma T}{b}\\ +c\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}\\[20.0pt] \begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}-4b\langle x^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}-4b\langle p^{2}\rangle\langle\{x,p\}\rangle\\ +2\gamma\end{array}&\begin{array}[]{c}-8b\langle x^{2}\rangle\langle p^{2}\rangle-2b\\ -2b\langle\{x,p\}\rangle^{2}\end{array}\\[20.0pt] \begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}4c\langle x^{2}\rangle\langle\{x,p\}\rangle\\ +2\gamma\langle\{x,p\}\rangle^{2}\end{array}&\begin{array}[]{c}8\gamma\langle p^{2}\rangle^{2}-\dfrac{8\gamma T}{b}\langle p^{2}\rangle\\ +4c\langle p^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}8c\langle x^{2}\rangle\langle p^{2}\rangle+2c+2c\langle\{x,p\}\rangle^{2}\\ +8\gamma\langle p^{2}\rangle\langle\{x,p\}\rangle-\dfrac{4\gamma T}{b}\langle\{x,p\}\rangle\end{array}\\[20.0pt] \begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}8c\langle x^{2}\rangle^{2}\\ +2b\\ +4\gamma\langle x^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}-8b\langle p^{2}\rangle^{2}-2c+4\gamma\langle p^{2}\rangle\langle\{x,p\}\rangle\\ +2c\langle\{x,p\}\rangle^{2}-\dfrac{4\gamma T}{b}\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}8\gamma\langle x^{2}\rangle\langle p^{2}\rangle-\dfrac{8\gamma T}{b}\langle x^{2}\rangle-2\gamma\\ +8c\langle x^{2}\rangle\langle\{x,p\}\rangle-8b\langle p^{2}\rangle\langle\{x,p\}\rangle\\ +2\gamma\langle\{x,p\}\rangle^{2}\end{array}\end{bmatrix}$ (45)
while the constant terms remain the same:
$D=\left[0,0,0,-2\frac{\gamma}{b},0\right].$ (46)
Interestingly the factorization between first order and second order terms still holds in this squeezed gaussian state. Solving for the expansion coefficients of the SLD we obtain three non-zero results (their lengthy expressions are given explicitly in appendix B and in the Mathematica scripts accompanying this manuscript [24]) that are functions of the non-vanishing expectation values in the system
$c^{(x^{2})}_{T}$ $=f(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle),$ (47)
$c^{(p^{2})}_{T}$ $=g(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle),$ (48)
$c^{(\{x,p\})}_{T}$ $=h(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle).$ (49)
We have checked explicitly that in the limit of
$\langle x^{2}\rangle\to\frac{a^{2}b}{4c},\langle p^{2}\rangle\to\frac{a^{2}}{4},\langle\{x,p\}\rangle\to 0,$ (50)
corresponding to the stationary late-time limit, we recover the same expressions as in eq. 39.
Let us explore the relative importance of the different operators contributing to the temperature SLD in the squeezed Gaussian state by solving the Ehrenfest equations of motion [7] associated with the CL master equation
$\frac{d}{dt}\langle x^{2}\rangle=\frac{1}{m}\langle px+xp\rangle,$ (51)
$\frac{d}{dt}\langle px+xp\rangle=\frac{2}{m}\langle p^{2}\rangle-2\langle m\omega x^{2}\rangle-2\gamma\langle px+xp\rangle,$ (52)
$\frac{d}{dt}\langle p^{2}\rangle=-m\omega\langle px+xp\rangle-4\gamma\langle p^{2}\rangle+4m\gamma T.$ (53)
Starting from the same initial conditions of $\langle x^{2}\rangle\langle p^{2}\rangle=1$ and $\langle\{x,p\}\rangle=1$, we choose two sets of evolution parameters that correspond to two qualitatively different scenarios. Making sure that $\gamma\ll 2\pi T$ and $\omega\ll 2\pi T$ we consider a probe particle within a bath at temperature $T/m=4$ and relaxation rate $\gamma/m=1/8$. In one case the confining potential is given by $\omega/m=4/3$ in the other by $\omega/m=2/3$.
Figure 2: Probe particle evolution in a squeezed Gaussian state at $\omega/m=4/3$ (left) and $\omega/m=2/3$ (right). Evolution of the expansion coefficients $c^{(i)}_{\gamma}$ of the symmetric logarithmic derivative for relaxation rate $c_{\gamma}^{xx}$ (green solid), $c_{\gamma}^{pp}$ (teal solid), $c_{\gamma}^{xp+px}$ (purple solid). All coefficients vanish in the late time stationary state.
In the potential with stronger curvature $\omega/m=4/3$, the particle is unable to diffuse over similarly large distances as in the the weaker potential $\omega/m=2/3$. Thus we expect differences in position and momentum spread to affect the composition of the SLD.
And indeed as shown in the fig. 1 that for $\omega/m=4/3$ (top row) momentum spread (teal solid) dominates over position spread (green solid), in contrast to the confining potential with $\omega/m=2/3$ (bottom row).
The corresponding expansion coefficients $c_{T}^{(i)}$ of the SLD (right column) appear to take on values, which weight the different operators roughly such that they contribute equally in the linear combination. I.e. in case that momentum spread dominates position spread by roughly a factor two (top row), the coefficient $c^{xx}_{T}$ is also roughly twice as large as $c^{pp}_{T}$. This reinforces the picture that it truly is the linear combination of position and momentum spread that constitutes the most sensitive probe of temperature. Similarly when position spread dominates, the coefficient $c^{pp}_{T}$ takes on a larger value than $c^{xx}_{T}$ (bottom row).
As all expansion coefficients in fig. 1 approach the correct equilibrium value in the late time limit, we are confident that our results are meaningful in the transient regime too.
IV.2 Relaxation rate metrology
Figure 3: Evolution of the ratio of expansion coefficients $c^{xx}/c^{pp}$ of the symmetric logarithmic derivative for temperature (green solid) and for the relaxation rate (purple solid) for the two cases of $\omega/m=4/3$ (left) and $\omega/m=2/3$ (right).
Having established that our formalism reproduces the correct SLD in equilibrium and produces predictions that are consistent with physical intuition, let us deploy it for a genuine non-equilibrium estimate, i.e. for constructing the SLD for the relaxation rate of the probe.
The CL master equation carries explicit factors of $\gamma$ in two terms, in contrast to temperature, which appears only once. But since the relaxation rate enter solely as prefactor to these terms, taking the derivative $\partial_{\gamma}(\partial_{t}\hat{\rho})$ leads to the same terms in the system matrix $M$ as in eq. 13 and the only difference lies in one additional term contributing to the constant contribution $D$
$D_{j}=\frac{T}{b}\left\langle\left[\hat{x},\left[\hat{x},\hat{A}_{j}\right]\right]\right\rangle+i\left\langle\left\{\hat{p},\left[\hat{A}_{j},\hat{x}\right]\right\}\right\rangle.$ (54)
Considering the same squeezed Gaussian state as before, we have as system matrix eq. 45 and for the constant terms we find
$D=\left[0,0,0,-2\frac{T}{b}+4\left\langle p^{2}\right\rangle,2\left\langle\{\hat{x},\hat{p}\}\right\rangle\right].$ (55)
The evolution of the coefficients is shown in fig. 2 for $\omega/m=4/3$ (left) and $\omega/m=2/3$ (right). As expected from the fact that in the steady state at late time no information about equilibration is retained, all coefficients approach zero over time. Again the ordering of the coefficients $c_{\gamma}^{xx}$ and $c_{\gamma}^{pp}$ follows the inverted hierarchy of the expectation values.
Figure 4: Evolution of the Quantum Fisher Information for temperature measurements (green solid) and for relaxation rate measurement (purple solid) for the two cases of $\omega/m=4/3$ (left) and $\omega/m=2/3$ (right). Note the qualitatively different behavior with ${\cal F}_{T}$ increasing towards late times, while ${\cal F}_{\gamma}$
Due to the simple dependence of the CL master equation on the environmental parameters $\gamma$ and $T$ we find that a similarly weighted linear combination of position and momentum spread constitutes an optimal estimator for both of these properties. Of course out of equilibrium, the mixed operator $\{x,p\}$ does contribute to the SLD but its expectation value decays towards zero. Figure 3 shows the explicit ratios of the coefficients $c^{xx}/c^{pp}$ for $\hat{\Lambda}_{T}$ (blue solid) and $\hat{\Lambda}_{\gamma}$ (orange solid) evaluated for $\omega/m=4/3$ (left) and $\omega/m=2/3$ (right). We find that for the parameters considered here, the ratios of these coefficients lie close together all throughout evolution towards the steady state.
Of course the explicit construction of the SLD is only part of the necessary steps to perform optimal temperature and relaxation rate metrology, as we also need to understand in which time intervals measurements are feasible. Intuitively, we expect that a fully thermally equilibrated probe will be most informative about the environment temperature, while it has lost all knowledge of the approach to equilibrium and thus becomes less and less informative of the relaxation rate.
And indeed, inspecting the Quantum Fisher Information ${\cal F}$ for temperature ${\cal F}_{T}$ and relaxation rate ${\cal F}_{\gamma}$, as shown in fig. 4 confirms our intuition. The QFI for temperature steadily increases until it has reached its maximum value at late times, while for the relaxation rate measurement the corresponding QFI decreases quickly over time, rendering the SLD irrelevant in the late time limit.
V Summary and Conclusion
We have introduced a general strategy for the explicit construction of optimally sensitive observables for quantum probes, based on the symmetric logarithmic derivative. By exploiting the symmetry of mixed derivatives guaranteed by Schwarz’s theorem, we relate the dynamical evolution governed by a master equation to the parameter sensitivity encoded in the SLD. Employing an operator basis expansion, our approach translates the identification of the SLD into the solution of a linear system for the associated expansion coefficients.
To validate the method, we applied it to the Caldeira–Leggett model in the high-temperature limit, where analytic expressions for the steady-state density matrix are available. In this setting, our construction successfully recovers the known SLD associated with temperature estimation in the late-time Gaussian state.
Crucially, our framework enables the determination of the SLD directly from the master equation, without requiring an explicit solution for the time-evolved density matrix. This novelty allows us to construct time-dependent expressions for the SLD even when the probe is out of equilibrium. We find that the structure of the optimal observable evolves over time and asymptotically approaches its thermal form in the long-time limit.
Access to the SLD in the transient regime opens new directions for quantum metrology beyond equilibrium. As an illustrative example, we derive the SLD associated with the environmental relaxation rate. Owing to the simple dependence of the Caldeira–Leggett dynamics on both relaxation rate and temperature, this observable turns out to be dominated by a linear combination of the position and momentum variance, which mirrors the structure of the temperature SLD.
With explicit SLD expressions at hand, we compute the corresponding quantum Fisher information, revealing that the thermalized probe indeed exhibits maximal sensitivity to temperature, whereas sensitivity to the relaxation rate is concentrated in the non-equilibrium regime. These findings highlight the potential of our approach to extend the practical applicability of the SLD in quantum metrology, particularly in regimes where the reduced state of the probe is not analytically accessible.
Acknowledgements.
AKR thanks Korea University for support through project K2503291 Ab-initio simulation of the real-time dynamics of non-relativistic fermions as well as project K2511131. VLP was supported by Xunta de Galicia project ED481A 2022/286, by European Research Council project ERC-2018-ADG-835105 YoctoLHC, by Xunta de Galicia (CIGUS Network of Research Centres), by European Union ERDF, by the Spanish Research State Agency under projects PID2020-119632GBI00 and PID2023-152762NB-I00, and project CEX2023-001318-M financed by MCIN/AEI/10.13039/501100011033
Appendix A Sketch of the derivation of the system matrix M
To obtain the expression eq. 13 of the matrix elements $M_{ij}$ in terms of expectation values, we start by taking the temperature derivative of the operator valued CL master equation eq. 11
$\partial_{T}\partial_{t}\hat{\rho}(t,T)$ $=-i[\hat{H},\partial_{T}\hat{\rho}]-i\gamma[\hat{x},\{\hat{p},\partial_{T}\hat{\rho}\}]$ (56)
$-2\gamma m[\hat{x},[\hat{x},\hat{\rho}]]-2\gamma mT[\hat{x},[\hat{x},\partial_{T}\hat{\rho}]],$
as well as the time derivative of the operator definition eq. 1 of the SLD, setting $\theta=T$, which yields
$\partial_{t}\partial_{T}\hat{\rho}(T)$ $=\frac{1}{2}\left(\hat{\Lambda}_{T}\partial_{t}\hat{\rho}(T)+\partial_{t}\hat{\rho}(T)\hat{\Lambda}_{T}\right)$ (57)
$-\partial_{t}\hat{\rho}(T)\langle\hat{\Lambda}_{T}\rangle-\hat{\rho}(T)\partial_{t}\langle\hat{\Lambda}_{T}\rangle,$
where $\partial_{t}\langle\hat{\Lambda}_{T}\rangle=\operatorname{Tr}[\partial_{t}\hat{\rho}(T)\hat{\Lambda}_{T}]$. Now, due to Schwartz’s theorem, we have eq. 7 and we can identify eq. 56 and eq. 57. Inserting eq. 11 and eq. 1 and excluding a time dependence of the SLD operator, we obtain an equation that contains no more explicit time or temperature derivatives.
To convert to expectation values, we make the operator expansion ansatz $\hat{\Lambda}_{T}=\sum_{i}c^{(i)}_{T}\hat{A}_{i}$ as in eq. 8 and take traces according to eq. 9. This allows us to obtain eq. 10 with $M_{ij}$ and $D_{j}$ as given by eq. 13 and eq. 14, respectively.
We can evaluate the individual entries of $M_{ij}$, for example $M_{x^{2}p^{2}}$ by inserting the corresponding expressions for $\hat{A}_{i}$ and $\hat{A}_{j}$ in eq. 13
$M_{x^{2}p^{2}}=$ (58)
$-\tfrac{1}{2}\left\langle\left\{\hat{x}^{2}\left[\hat{H},\hat{p}^{2}\right]\right\}\right\rangle-\tfrac{i}{2}\left\langle\left\{\hat{x}^{2},\left\{\hat{p},\left[\hat{p}^{2},\hat{x}\right]\right\}\right\}\right\rangle$
$-\tfrac{\gamma T}{2b}\left\langle\left\{\hat{x}^{2},\left[\hat{x},\left[\hat{x},\hat{p}^{2}\right]\right]\right\}\right\rangle+\tfrac{i\gamma}{2}\left\langle\left\{\hat{p},\left[\left\{\hat{x}^{2},\hat{p}^{2}\right\},\hat{x}\right]\right\}\right\rangle$
$+\tfrac{\gamma T}{2b}\left\langle\left[\hat{x},\left[\hat{x},\left\{\hat{x}^{2},\hat{p}^{2}\right\}\right]\right]\right\rangle-i\left\langle\hat{p}^{2}\right\rangle\left\langle\left[\hat{x}^{2},\hat{H}\right]\right\rangle$
$-i\gamma\left\langle\hat{p}^{2}\right\rangle\left\langle\left\{\hat{p},\left[\hat{x}^{2},\hat{x}\right]\right\}\right\rangle-\tfrac{\gamma T}{b}\left\langle\hat{p}^{2}\right\rangle\left\langle\left[\hat{x},\left[\hat{x},\hat{x}^{2}\right]\right]\right\rangle.$
Here we make use of know commutation relations
$[\hat{x},\hat{p}^{m}]=im\hat{p}^{m-1},\quad[\hat{x}^{n},\hat{p}]=in\hat{x}^{n-1},$
$[\hat{x},\{\hat{x},\hat{p}\}]=2i\hat{x},\quad[\hat{p},\{\hat{x},\hat{p}\}]=-2i\hat{p},$
$[\hat{x}^{2},\hat{p}^{2}]=2i\{\hat{x},\hat{p}\},$
as well as the Weyl correspondences collected in table 1 to obtain
$M_{x^{2}p^{2}}=-4b\langle W(xp^{3})\rangle+2\gamma+4b\langle p^{2}\rangle\langle W(xp)\rangle.$ (59)
This expression reproduces both the steady state result ($M_{x^{2}p^{2}}=2\gamma$) when using eq. 16 and eq. 17, as well as the squeezed Gaussian state result ($M_{x^{2}p^{2}}=-4b\langle p^{2}\rangle\langle\{\hat{x},\hat{p}\}\rangle+2\gamma$) when using eq. 43.
Appendix B Expansion coefficients of the temperature SLD in the squeezed gaussian state
For completeness and reproducibility, we provide below the explicit expressions for the non-vanishing expansion coefficients of the temperature SLD in the squeezed Gaussian state, introduced in the main text as eqs. 47 and 48 and eq. 49. The expressions are also provided as Mathematica scripts in the supplementary material [24].
$c^{(x^{2})}_{T}=f(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{f_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{f_{D}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (60)
$f_{N}$ $=18\left[\gamma\left\langle\left\{x,p\right\}\right\rangle\left(-4\left(9\left\langle p^{2}\right\rangle^{2}-1\right)\left\langle x^{2}\right\rangle\left(2T-\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle^{2}-18T^{2}+1\right)\right.$ (61)
$\left.+\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(18T\left\langle p^{2}\right\rangle-9\left\langle p^{2}\right\rangle^{2}-18T^{2}+1\right)\right.$
$\left.+\left(T-\left\langle p^{2}\right\rangle\right)\left(-9\gamma^{2}-4\left\langle x^{2}\right\rangle\left(\left(1-9\gamma^{2}\right)\left\langle p^{2}\right\rangle+9\left\langle p^{2}\right\rangle^{3}+18\gamma^{2}T\right)-9\left\langle p^{2}\right\rangle^{2}-1\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(T-\left\langle p^{2}\right\rangle\right)\left(-9\gamma^{2}-4\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+9\left\langle p^{2}\right\rangle^{2}\right)+\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(T-\left\langle p^{2}\right\rangle\right)\right]$
$f_{D}$ $=\gamma\left[18\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(-72T\left\langle p^{2}\right\rangle+72T^{2}-4\right)+9T\right)\right.$ (62)
$\left.-18\gamma\left\langle\left\{x,p\right\}\right\rangle\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(-9\left\langle p^{2}\right\rangle\left(4T\left\langle x^{2}\right\rangle+1\right)+\left(72T^{2}-4\right)\left\langle x^{2}\right\rangle+18T\right)-9\right)+9T-4\left\langle x^{2}\right\rangle\right)\right.$
$\left.-9\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(9\left(\gamma^{2}+\left\langle p^{2}\right\rangle^{2}\right)+4\left\langle x^{2}\right\rangle\left(2T-3\left\langle p^{2}\right\rangle\right)\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(9\left(9\gamma^{2}+2\right)+4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle\left(9\gamma^{2}+4\left\langle x^{2}\right\rangle\left(4T-3\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-2T\right)+1\right)+18T-4\left\langle x^{2}\right\rangle\right)\right)\right.$
$\left.-\left(16\left\langle x^{2}\right\rangle^{2}+9\right)\left(18T\left\langle p^{2}\right\rangle-9\left\langle p^{2}\right\rangle^{2}-1\right)\left(4\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+1\right)+162\gamma T\left\langle\left\{x,p\right\}\right\rangle{}^{5}-9\left\langle\left\{x,p\right\}\right\rangle{}^{6}\right]$
$c^{(p^{2})}_{T}=g(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{g_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{g_{D}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (63)
$g_{N}$ $=9\left[-9\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(8T-4\left\langle p^{2}\right\rangle\right)+1\right)-72\gamma\left\langle x^{2}\right\rangle\left\langle\left\{x,p\right\}\right\rangle\left(T-\left\langle p^{2}\right\rangle\right)\left(4T\left\langle x^{2}\right\rangle+1\right)\right.$ (64)
$\left.-\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(T-\left\langle p^{2}\right\rangle\right)\left(8\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle-2\left\langle x^{2}\right\rangle\right)+9\right)-\left(16\left\langle x^{2}\right\rangle^{2}+9\right)\left(T-\left\langle p^{2}\right\rangle\right)\left(4\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+1\right)-9\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{5}\right]$
$g_{D}$ $=2\gamma\left[18\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(-72T\left\langle p^{2}\right\rangle+72T^{2}-4\right)+9T\right)\right.$
$\left.-18\gamma\left\langle\left\{x,p\right\}\right\rangle\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(-9\left\langle p^{2}\right\rangle\left(4T\left\langle x^{2}\right\rangle+1\right)+\left(72T^{2}-4\right)\left\langle x^{2}\right\rangle+18T\right)-9\right)+9T-4\left\langle x^{2}\right\rangle\right)\right.$
$\left.-9\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(9\left(\gamma^{2}+\left\langle p^{2}\right\rangle^{2}\right)+4\left\langle x^{2}\right\rangle\left(2T-3\left\langle p^{2}\right\rangle\right)\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(9\left(9\gamma^{2}+2\right)+4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle\left(9\gamma^{2}+4\left\langle x^{2}\right\rangle\left(4T-3\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-2T\right)+1\right)+18T-4\left\langle x^{2}\right\rangle\right)\right)\right.$
$\left.-\left(16\left\langle x^{2}\right\rangle^{2}+9\right)\left(18T\left\langle p^{2}\right\rangle-9\left\langle p^{2}\right\rangle^{2}-1\right)\left(4\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+1\right)+162\gamma T\left\langle\left\{x,p\right\}\right\rangle{}^{5}-9\left\langle\left\{x,p\right\}\right\rangle{}^{6}\right]$
$c^{(\{x,p\})}_{T}=h(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{h_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{h_{D}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (65)
$h_{N}$ $=9\left[4\gamma\left\langle x^{2}\right\rangle\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(-9\left\langle p^{2}\right\rangle^{2}+18T^{2}-1\right)-\gamma\left(16\left\langle x^{2}\right\rangle^{2}+9\right)\left(T-\left\langle p^{2}\right\rangle\right)\right.$ (66)
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(4\left\langle x^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-T\right)-9\gamma^{2}\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle\left(T-\left\langle p^{2}\right\rangle\right)\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle+4\left\langle x^{2}\right\rangle\right)+9\right)+4\left(1-9\gamma^{2}\right)\left\langle x^{2}\right\rangle\right)+9\gamma\left\langle p^{2}\right\rangle\left\langle\left\{x,p\right\}\right\rangle{}^{4}\right]$
$h_{D}$ $=\gamma\left[18\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(-72T\left\langle p^{2}\right\rangle+72T^{2}-4\right)+9T\right)\right.$ (67)
$\left.-18\gamma\left\langle\left\{x,p\right\}\right\rangle\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(-9\left\langle p^{2}\right\rangle\left(4T\left\langle x^{2}\right\rangle+1\right)+\left(72T^{2}-4\right)\left\langle x^{2}\right\rangle+18T\right)-9\right)+9T-4\left\langle x^{2}\right\rangle\right)\right.$
$\left.-9\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(9\left(\gamma^{2}+\left\langle p^{2}\right\rangle^{2}\right)+4\left\langle x^{2}\right\rangle\left(2T-3\left\langle p^{2}\right\rangle\right)\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(9\left(9\gamma^{2}+2\right)+4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle\left(9\gamma^{2}+4\left\langle x^{2}\right\rangle\left(4T-3\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-2T\right)+1\right)+18T-4\left\langle x^{2}\right\rangle\right)\right)\right.$
$\left.-\left(16\left\langle x^{2}\right\rangle^{2}+9\right)\left(18T\left\langle p^{2}\right\rangle-9\left\langle p^{2}\right\rangle^{2}-1\right)\left(4\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+1\right)+162\gamma T\left\langle\left\{x,p\right\}\right\rangle{}^{5}-9\left\langle\left\{x,p\right\}\right\rangle{}^{6}\right]$